An old stamp machine at the post office is very cantankerous and will only take nickels and quarters. If there are 45 coins totaling $10.25 when it is emptied st noon, how many nickels are there?
Answer by nerdybill(7384)

An old stamp machine at the post office is very cantankerous and will only take nickels and quarters. If there are 45 coins totaling $10.25 when it is emptied st noon, how many nickels are there?
.
Let n = number of nickels
then
45-n = number of quarters
.
.05n + .25(45-n) = 10.25
.05n + 11.25 - .25n = 10.25
11.25 - .20n = 10.25
-.20n = -1
n = (-1)/(-.20)
n = 5 (number of nickels)